[Hallc_running] [New Logentry] RC Report - 8/28/22

arunts at jlab.org arunts at jlab.org
Sun Aug 28 11:45:02 EDT 2022

Logentry Text:
We took data on dummy target at 40 uA and then moved to LD2 and took more data at 14 uA for tracking efficiency studies by Nathan (https://logbooks.jlab.org/entry/4034888). Based on the results, 12 uA seemed like a comfortable spot to continue taking data at. We had a very productive session until the test lab power outage at 5 am this morning (https://logbooks.jlab.org/entry/4034955) during which the HV for the GEMS in the prototype detector were turned off. Test Lab power and site chilled water power has been restored at ~10:30 am and we should expect beam in 1-2 hours. The plan is to move on to the next setting during the beam restoration process, and take data at a beam current where:

1. 700 kHz < pTrig rates < 800 kHz (based on what we found from the results from the previous setting)
2.  pTrig5 (coincidence) rates  ~ 3 kHz
3. deadtime is reasonable

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071

HMS NMR Dipole limits adjusted to avoid constant alarms: https://logbooks.jlab.org/entry/4034772, https://logbooks.jlab.org/entry/4034770
Rates at 14 uA: https://logbooks.jlab.org/entry/4034686
Brief SHMS interlock problems: https://logbooks.jlab.org/entry/4034937
HMS Q1 interlock#2 which cleared after two resets: https://logbooks.jlab.org/entry/4034877


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4034998
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220828/2c9122a1/attachment.html>

More information about the Hallc_running mailing list